A slender rod 40 in long is bent so as to form a right triangle. If the segments are 8 in and 32 in long, find the centroid.


1
Expert's answer
2021-07-26T15:39:58-0400

The centroid of a triangle is the point of intersection of its medians.

The centroid of a triangle divides each median in the ratio 2:1.

If the vertical segment is 32 in up and horisontal segment is 8 in right from the vertex of the right angle, then the first median is "y=-2x+16" and the second median is "y=-8x+32."

Intersection


"-2x+16=-8x+32"




"6x=16"

"x=\dfrac{8}{3}"

"y=-2(\dfrac{8}{3})+16=\dfrac{32}{3}"

The centroid is (32/3) in up and (8/3) in right from the vertex of the right angle,